Annales Academiæ Scientiarum Fennicæ Mathematica Volumen 31, 2006, 71–95

REMOVABLE SINGULARITIES FOR BOUNDED p-HARMONIC AND QUASI(SUPER)HARMONIC FUNCTIONS ON METRIC SPACES

Anders Björn

Linköpings universitet, Department of Mathematics SE-581 83 Linköping, Sweden; anbjo@mai.liu.se

Abstract. We study removable singularities for bounded p-harmonic functions in complete doubling metric spaces supporting a Poincaré inequality. We show that a relatively closed set E of an arbitrary open set Ω is removable if it has zero capacity. Moreover, the extensions are unique.

We give several general results showing nonremovability and giving us characterizations of removable sets as sets of capacity zero in various situations.

We also provide examples of removable sets with positive capacity. Such examples can have some unexpected behaviour: e.g., E may disconnect Ω; the extensions may not be unique; and there exists a nonremovable union of two compact disjoint removable sets.

Similar results for superharmonic, quasiharmonic and quasisuperharmonic functions are also given.

1. Introduction

(Relatively closed) sets of zero capacity are removable for bounded p-har-monic and bounded superharmonic functions defined in an open subset of weighted R^{n}, see Theorems 7.35 and 7.36 in Heinonen–Kilpeläinen–Martio [12]. They also show, p. 143 in [12], that compact sets are removable if and only if they have capacity zero. (See Serrin [28], [29] and Maz’ya [27] for earlier proofs of this fact for unweighted R^{n}.)

In Björn–Björn–Shanmugalingam [10, Section 8], the study of removable sets for bounded p-harmonic and superharmonic functions was extended to bounded domains Ω (with complement of positive capacity) in metric spaces. Removability was shown for sets of capacity zero.

The sharpness of the removability results in [10] was illustrated by showing nonremovability for a large class of sets with positive capacity. However, they were not able to show nonremovability for all sets with positive capacity, as pointed out on p. 419 in [10]. The reason for this turns out to be quite natural since we here present a simple counterexample, see Section 9.

2000 Mathematics Subject Classification: Primary 31C45; Secondary 35J65, 49J27, 49N60.

Let Ω be an open set in our metric space and E ⊂ Ω be relatively closed with zero capacity. In addition to showing that E is removable for bounded pharmonic functions and superharmonic functions bounded from below, we also show that E is removable for p-harmonic and superharmonic functions belonging to the Newtonian–Sobolev space N^{1,p}(Ω \ E). We also obtain the corresponding results for quasi(super)harmonic functions. (See Section 3 for the definition of quasi(super)harmonic functions and Section 6 for the precise statements of the removability results.)

It should be mentioned that for unweighted R^{n} Tolksdorf [33, Theorem 1.5], proved that compact sets of capacity zero are removable for bounded quasisuperminimizers (which is equivalent to removability for bounded quasisuperharmonic functions, by Proposition 6.6 below).

In Section 7 we extend the nonremovability results of [10] to quasi(super)- harmonic functions, as well as providing some more nonremovability results. This leads to characterizations of removable sets as those of capacity zero in several cases, see Section 8.

In Section 9 we give examples of sets with positive capacity that are removable for bounded p-harmonic and bounded Q-quasiharmonic functions. When a set of capacity zero is removed the extensions are unique. However, when a set of positive capacity is removable it is possible that a given p-harmonic function has several extensions. We provide two examples of such nonunique removability in Section 10. (A necessary requirement for nonunique removability is of course nonunique continuation of p-harmonic functions; see Martio [26] and Björn–Björn– Shanmugalingam [10, pp. 426–427].) We also give examples of removable sets E which disconnect Ω.

In Section 11 we point out that removable singularities with positive capacity have some unexpected behaviour, e.g. it is possible to have a nonremovable union of two compact disjoint removable singularities.

For more on quasiminimizers and their importance see the introductions in Kinnunen–Martio [23] and A. Björn [4]. An application of the removability results for quasiharmonic functions will be given in the forthcoming paper A. Björn [7].

For examples of complete metric spaces equipped with a doubling measure supporting a Poincaré inequality, see, e.g., A. Björn [3], [5].

Acknowledgement. The author is supported by the Swedish Research Council and Gustaf Sigurd Magnuson’s fund of the Royal Swedish Academy of Sciences. This research started while the author was visiting the Department of Mathematical Analysis at the Charles University in Prague during the autumn 2003.

2. Notation and preliminaries

Removable singularities for bounded p-harmonic functions 73

We assume throughout the paper that X = (X, d, μ) is a complete metric space endowed with a metric d and a doubling measure μ, i.e. there exists a constant C > 0 such that for all balls B = B(x_{0}, r) := {x ∈ X : d(x, x_{0}) < r} in X (we make the convention that balls are nonempty and open),

0 < μ(2B) ≤ Cμ(B) < ∞,

where λB = B(x_{0}, λr). We emphasize that the σ -algebra on which μ is defined is obtained by completion of the Borel σ -algebra. We also assume that 1 < p < ∞ and that Ω ⊂ X is a nonempty open set. (At the end of this section we make some further assumptions assumed in the rest of the paper.)

Note that some authors assume that X is proper (i.e. closed bounded sets are compact) rather than complete, but, since μ is doubling, X is proper if and only if it is complete.

A curve is a continuous mapping from an interval. We will in addition, throughout the paper, assume that every curve is nonconstant, compact and rectifiable. A curve can thus be parameterized by its arc length ds.

Definition 2.1. A nonnegative Borel function g on X is an upper gradient of an extended real-valued function f on X if for all curves γ: [0, l_{γ}] → X,

(2.1)  |f(γ(0)) − f(γ(l_{γ}))| ≤ ∫_{γ} g ds

whenever both f(γ(0)) and f(γ(l_{γ})) are finite, and ∫_{γ} g ds = ∞ otherwise. If g is a nonnegative measurable function on X and if (2.1) holds for p-almost every curve, then g is a p-weak upper gradient of f.

By saying that (2.1) holds for p-almost every curve we mean that it fails only for a curve family with zero p-modulus, see Definition 2.1 in Shanmugalingam [30]. It is implicitly assumed that ∫_{γ} g ds is defined (with a value in [0, ∞]) for p-almost every curve.

If g ∈ L^{p}(X) is a p-weak upper gradient of f, then one can find a sequence {g_{j}}_{j=1}^{∞} of upper gradients of f such that g_{j} → g in L^{p}(X), see Lemma 2.4 in Koskela–MacManus [25].

If f has an upper gradient in L^{p}(X), then it has a minimal p-weak upper gradient g_{f} ∈ L^{p}(X) in the sense that g_{f} ≤ g μ-a.e. for every p-weak upper gradient g ∈ L^{p}(X) of f, see Corollary 3.7 in Shanmugalingam [31].

If f, h ∈ N^{1,p}(X), then g_{f} = g_{h} μ-a.e. in {x ∈ X : f(x) = h(x)}, in particular g_{min{f,c}} = g_{f}χ_{f≠c} for c ∈ R. For these and other facts on p-weak upper gradients, see, e.g., Björn–Björn [8, Section 3].

Definition 2.2. We say that X supports a weak (1, q)-Poincaré inequality if there exist constants C > 0 and λ ≥ 1 such that for all balls B ⊂ X, all measurable functions f on X and all upper gradients g of f,

(2.2)  \overline{∫}_{B} |f − f_{B}| dμ ≤ C diam(B)(\overline{∫}_{λB} g^{q} dμ)^{1/q},

where f_{B} := \overline{∫}_{B} f dμ := μ(B)^{−1} ∫_{B} f dμ.

In the definition of Poincaré inequality we can equivalently assume that g is a q -weak upper gradient—see the comments above. It is also equivalent to require that (2.2) holds for all f ∈ Lip_{c}(X) and all upper gradients g ∈ Lip_{c}(X) of f, see Keith [18, Theorem 2]. We say that E ⋐ A if \overline{E} is a compact subset of A, and let Lip_{c}(A) = {f ∈ Lip(A) : supp f ⋐ A}.

Following Shanmugalingam [30], we define a version of Sobolev spaces on the metric space X.

Definition 2.3. Whenever u ∈ L^{p}(X), let

∥u∥_{N^{1,p}(X)} = (∫_{X} |u|^{p} dμ + inf_{g} ∫_{X} g^{p} dμ)^{1/p},

where the infimum is taken over all upper gradients of u. The Newtonian space on X is the quotient space

N^{1,p}(X) = {u : ∥u∥_{N^{1,p}(X)} < ∞}/∼,

where u ∼ v if and only if ∥u − v∥_{N^{1,p}(X)} = 0.

The space N^{1,p}(X) is a Banach space and a lattice, see Shanmugalingam [30].

Definition 2.4. The capacity of a set E ⊂ X is the number

C_{p}(E) = inf ∥u∥_{N^{1,p}(X)}^{p},

where the infimum is taken over all u ∈ N^{1,p}(X) such that u = 1 on E.

The capacity is countably subadditive. For this and other properties as well as equivalent definitions of the capacity we refer to Kilpeläinen–Kinnunen–Martio [20] and Kinnunen–Martio [21], [22].

We say that a property regarding points in X holds quasieverywhere (q.e.) if the set of points for which the property does not hold has capacity zero. The capacity is the correct gauge for distinguishing between two Newtonian functions. If u ∈ N^{1,p}(X), then u ∼ v if and only if u = v q.e. Moreover, Corollary 3.3 in Shanmugalingam [30] shows that if u, v ∈ N^{1,p}(X) and u = v μ-a.e., then u ∼ v.

If X supports a weak (1, p)-Poincaré inequality, then Lipschitz functions are dense in N^{1,p}(X) and the functions in N^{1,p}(X) are quasicontinuous, see Remark 4.4 in [30]. This means that in the Euclidean setting, N^{1,p}(R^{n}) is the refined Sobolev space as defined on p. 96 of Heinonen–Kilpeläinen–Martio [12].

To be able to compare the boundary values of Newtonian functions we need a Newtonian space with zero boundary values. We let for a measurable set E ⊂ X,

N_{0}^{1,p}(E) = {f|_{E} : f ∈ N^{1,p}(X) and f = 0 on X \ E}.

One can replace the assumption “f = 0 on X \ E ” with “f = 0 q.e. on X \ E ” without changing the obtained space N_{0}^{1,p}(E). Note that if C_{p}(X \ E) = 0, then N_{0}^{1,p}(E) = N^{1,p}(E).

We say that f ∈ N_{loc}^{1,p}(Ω) if f ∈ N^{1,p}(Ω′) for every open Ω′ ⋐ Ω.

By a continuous function we always mean a real-valued continuous function, whereas a semicontinuous function is allowed to be extended real-valued, i.e. to take values in the extended real line R := [−∞, ∞]. We let f_{+} = max{f, 0} and f_{−} = max{−f, 0}.

In addition to the assumptions made in the beginning of this section, from now on we assume that X supports a weak (1, p)-Poincaré inequality. By Keith– Zhong [19] it follows that X supports a weak (1, q)-Poincaré inequality for some q ∈ [1, p), which was earlier a standard assumption. Throughout the paper we also let Q ≥ 1 be a real number.

3. Quasi(super)harmonic functions

We follow Kinnunen–Martio [23, Section 3], making the following definition.

Definition 3.1. A function u ∈ N_{loc}^{1,p}(Ω) is a Q-quasiminimizer in Ω if for all open Ω′ ⋐ Ω and all φ ∈ N_{0}^{1,p}(Ω′) we have

(3.1)  ∫_{Ω′} g_{u}^{p} dμ ≤ Q ∫_{Ω′} g_{u+φ}^{p} dμ.

A function u ∈ N_{loc}^{1,p}(Ω) is a Q-quasisuperminimizer in Ω if (3.1) holds for all nonnegative functions φ ∈ N_{0}^{1,p}(Ω′), and a Q-quasisubminimizer in Ω if (3.1) holds for all nonpositive functions φ ∈ N_{0}^{1,p}(Ω′).

A function is a Q-quasiminimizer in Ω if and only if it is both a Q-quasisub-minimizer and a Q-quasisuperminimizer in Ω.

We will need some characterizations of Q-quasisuperminimizers. The following result was proved in A. Björn [4].

Proposition 3.2. Let u ∈ N_{loc}^{1,p}(Ω). Then the following are equivalent:

(a) The function u is a Q-quasisuperminimizer in Ω.

(b) For all nonnegative φ ∈ Lip_{c}(Ω) we have

∫_{φ≠0} g_{u}^{p} dμ ≤ Q ∫_{φ≠0} g_{u+φ}^{p} dμ.

(c) For all nonnegative φ ∈ N_{0}^{1,p}(Ω) we have

∫_{φ≠0} g_{u}^{p} dμ ≤ Q ∫_{φ≠0} g_{u+φ}^{p} dμ.

If we omit “super” from (a) and “nonnegative” from (b) and (c) we have a corresponding characterization for Q-quasiminimizers.

By Proposition 3.8 and Corollary 5.5 in Kinnunen–Shanmugalingam [24], a Q-quasiminimizer can be modified on a set of capacity zero so that it becomes locally Hölder continuous in Ω. A Q-quasiharmonic function is a continuous Q-quasiminimizer.

Kinnunen–Martio [23, Theorem 5.3], showed that if u is a Q-quasisupermini-mizer in Ω, then its lower semicontinuous regularization

u^{∗}(x) := ess lim inf_{y→x} u(y)

is also a Q-quasisuperminimizer in Ω in the same equivalence class as u in N_{loc}^{1,p}(Ω). Furthermore, u^{∗} is Q-quasisuperharmonic in Ω.

Definition 3.3. A function u: Ω → (−∞, ∞] is Q-quasisuperharmonic in Ω if u is not identically ∞ in any component of Ω, and there is a sequence of open sets Ω_{j} and Q-quasisuperminimizers v_{j}: Ω_{j} → (−∞, ∞] such that

(i) Ω_{j} ⋐ Ω_{j+1} ; (ii) Ω = ⋃_{j=1}^{∞} Ω_{j}; (iii) v_{j} ≤ v_{j+1} in Ω_{j}; (iv) u = lim_{j→∞} v_{j}^{∗} in Ω,

where v_{j}^{∗} is the lower semicontinuous regularization of v_{j}.

This definition is due to Kinnunen–Martio [23, Definition 7.1]. The following characterization, Theorem 7.10 in [23], is often useful. (Note that there are misprints in Definition 7.1 and Theorem 7.10 in [23]—which have been corrected here.)

Theorem 3.4. A function u: Ω → (−∞, ∞] is Q-quasisuperharmonic in Ω if u is not identically ∞ in any component of Ω, u is lower semicontinuously regularized, and min{u, k} is a Q-quasisuperminimizer in Ω for every k ∈ R.

If u_{j} is a Q_{j} -quasisuperminimizer in Ω, j = 1, 2, then, by Corollary 3.8 in [23], min{u_{1}, u_{2}} is a min{Q_{1} + Q_{2}, Q_{1}Q_{2}}-quasisuperminimizer in Ω; there is also a corresponding result for quasisuperharmonic functions, see Theorem 7.6 in [23]. We will use these facts mainly with u_{2} constant.

By Lemma 5.2 in [23], a quasisuperharmonic function u in Ω obeys the minimum principle: If u(x) = inf_{Ω} u for some x ∈ Ω, then u is constant in the component of Ω containing x.

Corollary 3.5. Assume that X is bounded and that u is a quasisuperharmonic function on X. Then u is constant.

It may be worth to observe that even if our removability results later in the paper will show that if C_{p}(X \ Ω) = 0, then any quasisuperharmonic function on Ω which is either bounded from below or in N^{1,p}(Ω) is constant, there may still exist unbounded quasisuperharmonic functions on Ω. Let, e.g., X = \overline{B(0, 1)} in (unweighted) R^{3} , p = 2, Ω = X \ {0} and u(x) = −|x|^{−1} . Then it is not difficult to show that u is 2-superharmonic in Ω. (Note however that by, e.g, the maximum principle u is not 2-harmonic in Ω.)

Proof. Since X is compact and u lower semicontinuous, there is x such that u(x) = inf_{X} u. As X is connected (which follows from the Poincaré inequality), the minimum principle gives that u is constant.

The following boundary minimum principle was proved in A. Björn [4].

Lemma 3.6. Assume that Ω is bounded, C_{p}(X \ Ω) > 0 and f ∈ C(∂Ω) ∩ N_{0}^{1,p}(\overline{Ω}). Let u be a quasisuperharmonic function in Ω satisfying u−f ∈ N_{0}^{1,p}(Ω). Then inf_{Ω} u ≥ inf_{∂Ω} f.

We will need the following convergence result.

Proposition 3.7. Let {u_{j}}_{j=1}^{∞} be a nondecreasing sequence of Q-quasisuper-harmonic functions in Ω and let u = lim_{j→∞} u_{j}. Then either u is identically ∞ in some component of Ω or u is Q-quasisuperharmonic in Ω.

Proof. Let k ∈ R, w_{k} = min{u, k} and w_{k,j} = min{u_{j}, k}. By Theorem 3.4, w_{k,j} is a lower semicontinuously regularized Q-quasisuperminimizer in Ω.

Let further Ω_{1} ⋐ Ω_{2} ⋐ ⋯, be open and such that Ω = ⋃_{j=1}^{∞} Ω_{j}. Using w_{k} and w_{k,j} in the place of u and v_{j}, respectively, in Definition 3.3 we see that w_{k} is Q-quasisuperharmonic in Ω.

Hence w_{k} is lower semicontinuously regularized and it follows easily that also u is lower semicontinuously regularized. Thus using Theorem 3.4 we see that u is Q-quasisuperharmonic in Ω, or identically ∞ in some component.

The corresponding result for locally uniformly convergent sequences can be obtained as a corollary.

Corollary 3.8. Let {u_{j}}_{j=1}^{∞} be a sequence of Q-quasisuperharmonic functions in Ω which converges locally uniformly and let u = lim_{j→∞} u_{j}. Then u is Q-quasisuperharmonic in Ω.

Proof. Let Ω′ ⋐ Ω be open. Then u_{j} → u uniformly in Ω′, and by choosing a subsequence if necessary we may assume that sup_{Ω′} |u_{j} − u| < 2^{−j}. Let now v_{j} = u_{j} − 3 · 2^{−j}. It is easy to check that v_{j} ≤ v_{j+1} in Ω′. Moreover v_{j} → u in Ω′, and thus by Proposition 3.7, u is Q-quasisuperharmonic in Ω′. Since Ω′ was an arbitrary compactly contained open subset of Ω we conclude that u is Q-quasisuperharmonic in Ω.

4. A Caccioppoli type estimate for quasisubminimizers

We will need a Caccioppoli type estimate for quasisubminimizers. This estimate was proved for unweighted R^{n} by Tolksdorf [33, Theorem 1.4]. The proof given here is an easy adaptation of Tolksdorf’s proof to metric spaces.

Theorem 4.1. Let u ≥ 0 be a Q-quasisubminimizer in Ω. Then for all nonnegative η ∈ Lip_{c}(Ω),

∫_{Ω} g_{u}^{p}η^{p} dμ ≤ c ∫_{Ω} u^{p}g_{η}^{p} dμ,

where c only depends on p and Q.

Proof. After multiplying η with a constant we may assume that 0 ≤ η ≤ 1. Let φ = ηu ∈ N_{0}^{1,p}(Ω). By the quasisubminimizing property of u we have

∫_{η=1} g_{u}^{p} dμ ≤ ∫_{φ>0} g_{u}^{p} dμ ≤ Q ∫_{φ>0} g_{u−φ}^{p} dμ ≤ Q ∫_{η>0} g_{u(1−η)}^{p} dμ.

Using that g_{u(1−η)} ≤ ug_{1−η} +(1−η)g_{u} = ug_{η} +(1−η)g_{u} we get with Q_{2} = 2^{p−1}Q,

\begin{aligned}∫_{η=1} g_{u}^{p} dμ & ≤ Q_{2} ∫_{η>0} u^{p}g_{η}^{p} dμ + Q_{2} ∫_{η>0}(1 − η)^{p}g_{u}^{p} dμ \\ & ≤ Q_{2} ∫_{Ω} u^{p}g_{η}^{p} dμ + Q_{2} ∫_{0<η<1} g_{u}^{p} dμ. \end{aligned}

By adding Q_{2} times the left-hand side to both sides, and then dividing by Q_{2} +1, we obtain

(4.1)  ∫_{η=1} g_{u}^{p} dμ ≤ θ ∫_{Ω} u^{p}g_{η}^{p} dμ + θ ∫_{η>0} g_{u}^{p} dμ,

where θ = Q_{2}/(Q_{2} + 1) < 1.

Next let τ be such that 0 < τ^{−p}θ < τ^{−2p}θ < 1 (note that it follows that 0 < τ < 1), η_{j} = (min{η, τ^{j}} − τ^{j+1})_{+} and η̂_{j} = η_{j}/(τ^{j} − τ^{j+1}). Applying (4.1) to η̂_{j} gives

\begin{aligned}∫_{Ω} g_{u}^{p}η̂_{j−1}^{p} dμ & ≤ ∫_{η̂_{j}=1} g_{u}^{p} dμ ≤ θ ∫_{Ω} u^{p}gp ; η̂_{j} dμ + θ ∫_{η̂_{j}>0} g_{u}^{p} dμ \\ & ≤ θ ∫_{Ω} u^{p}gp ; η̂_{j} dμ + θ ∫_{Ω} g_{u}^{p}η̂_{j+1}^{p} dμ \end{aligned}

from which it follows that

∫_{Ω} g_{u}^{p}η_{j−1}^{p} dμ ≤ τ^{−p}θ ∫_{Ω} u^{p}g_{η_{j}}^{p} dμ + τ^{−2p}θ ∫_{Ω} g_{u}^{p}η_{j+1}^{p} dμ.

Using that g_{η_{j}} ≤ g_{η} and iterating gives us

(4.2)  ∫_{η=1} g_{u}^{p} dμ ≤ (1 − τ)^{−p} ∫_{Ω} g_{u}^{p}η_{0}^{p} dμ ≤ Q_{3} ∫_{Ω} u^{p}g_{η}^{p} dμ,

where Q_{3} = (1 − τ)^{−p}τ^{−p}θ(1 − τ^{−2p}θ)^{−1} . Using (4.2) with η replaced by ψ = min{(4η − 1)_{+}, 1} gives

∫_{η≥1/2} g_{u}^{p} dμ = ∫_{ψ=1} g_{u}^{p} dμ ≤ Q_{3} ∫_{Ω} u^{p}g_{ψ}^{p} dμ ≤ 4^{p}Q_{3} ∫_{1/4<η<1/2} u^{p}g_{η}^{p} dμ.

Finally, let ψ_{j} = min{2^{j}η, 1}. Then we get

\begin{aligned}∫_{Ω} g_{u}^{p}η^{p} dμ & ≤ ∑_{j=0}^{∞} 2^{−jp} ∫_{ψ_{j}≥1/2} g_{u}^{p} dμ \\ & ≤ 4^{p}Q_{3} ∑_{j=0}^{∞} 2^{−jp} ∫_{1/4<ψ_{j}<1/2} u^{p}g_{ψ_{j}}^{p} dμ ≤ 4^{p}Q_{3} ∫_{Ω} u^{p}g_{η}^{p} dμ. \end{aligned}

5. p-harmonic functions

If Q = 1, “quasi” is omitted from the notation and a function is, e.g., pharmonic if it is 1-quasiharmonic, and superharmonic if it is 1-quasisuperhar-monic. For equivalent definitions and characterizations of superharmonic functions, see A. Björn [5].

We need some results for p-harmonic functions, see, e.g., Shanmugalingam [31] or Björn–Björn–Shanmugalingam [9]. Assume throughout this section that Ω is bounded and C_{p}(X \ Ω) > 0.

If f ∈ N^{1,p}(X), then there is a unique solution to the Dirichlet problem with boundary values f, i.e. there is a unique function Hf = H_{Ω}f such that Hf = f in X \ Ω and Hf is p-harmonic in Ω.

Lemma 5.1. If f_{1}, f_{2} ∈ N^{1,p}(X) and f_{1} ≤ f_{2} q.e. on ∂Ω, then Hf_{1} ≤ Hf_{2} in Ω.

It follows that for f ∈ N^{1,p}(X), (Hf)|_{Ω} only depends on f|_{∂Ω} . A Lipschitz function f on ∂Ω can be extended to a function f̃ ∈ Lip_{c}(X) such that f = f̃ on ∂Ω. As (Hf̃)|_{Ω} does not depend on the choice of extension, we define Hf := (Hf̃)|_{Ω} .

Definition 5.2. A point x_{0} ∈ ∂Ω is regular if

lim_{Ω∋y→x_{0}} Hf(y) = f(x_{0}) for all f ∈ Lip(∂Ω).

If x_{0} ∈ ∂Ω is not regular, then it is irregular.

For equivalent characterizations of regular boundary points see Björn–Björn [8, Theorem 6.2]. Recall the following result from Björn–Björn–Shanmugalingam [9, Theorem 3.9].

Theorem 5.3 (The Kellogg property). The set of all irregular points on ∂Ω has capacity zero.

6. Removability

From now on let E ⊊ Ω be relatively closed and such that no component of Ω is completely contained in E.

Definition 6.1. The set E is (uniquely) removable for bounded Q-quasi- (super)harmonic functions in Ω \ E if every bounded Q-quasi(super)harmonic function u in Ω \ E has a (unique) bounded Q-quasi(super)harmonic extension to Ω.

Similarly, E is (uniquely) removable for Q-quasi(super)harmonic functions in N^{1,p}(Ω \ E) if every Q-quasi(super)harmonic function u ∈ N^{1,p}(Ω \ E) has a (unique) Q-quasi(super)harmonic extension in N^{1,p}(Ω).

Here, when we, e.g., talk about a Q-quasiharmonic function u ∈ N^{1,p}(Ω), we mean that it is Q-quasiharmonic in Ω.

In view of Theorems 6.2 and 6.3 below one can get the feeling that it is not essential to stress the set Ω \ E in the definition, and that the notation could therefore be simplified. However, in Section 11, we see in (a′) that for removable singularities of positive capacity it is essential to always stress on which set we discuss removability.

In this section we show that sets of capacity zero are removable for Q-quasi- (super)harmonic functions that are either bounded or in N^{1,p}. Moreover the extensions are always unique in these cases.

In Björn–Björn–Shanmugalingam [10, Propositions 8.2 and 8.3], the following two results were shown under the additional assumption that Q = 1, u is bounded and Ω is a bounded domain with C_{p}(X \ Ω) > 0. (The uniqueness was given as a comment in the text.) In the proof of Proposition 8.3 (which was also used to derive Proposition 8.2) in [10] it was not explicitly shown that u ∈ N_{loc}^{1,p}(Ω). How this can be observed is pointed out in the proof of Theorem 6.3 below.

Theorem 6.2. Assume that C_{p}(E) = 0. Let u be a Q-quasiharmonic function in Ω \ E. Assume that one of the following conditions hold:

(a) u ∈ N^{1,p}(Ω \ E);

(b) u is the restriction of a function in N_{loc}^{1,p}(Ω);

(c) u ∈ N^{1,p}(B \ E) for every ball B ⋐ Ω;

(d) u is bounded. Then u has a unique quasisuperharmonic extension U to Ω. Moreover, U is Q-quasiharmonic in Ω, U is bounded if u is bounded, and U ∈ N^{1,p}(Ω) if u ∈ N^{1,p}(Ω \ E).

By saying that u is quasisuperharmonic we, of course, mean that there is some Q′ such that u is Q′ -quasisuperharmonic.

Theorem 6.3. Assume that C_{p}(E) = 0. Let u be a Q-quasisuperharmonic function in Ω \ E. Assume that one of the following conditions hold:

(a) u ∈ N^{1,p}(Ω \ E);

(b) u is the restriction of a function in N_{loc}^{1,p}(Ω);

(c) u ∈ N^{1,p}(B \ E) for every ball B ⋐ Ω;

(d) u is bounded from below.

Then u has a unique quasisuperharmonic extension U to Ω. Moreover, U is Q-quasisuperharmonic in Ω, U is bounded if u is bounded, and U ∈ N^{1,p}(Ω) if u ∈ N^{1,p}(Ω \ E).

Note that to find the extension U we only need to find the unique semicontinuously regularized extension of u to Ω; U(x) = ess lim inf_{Ω\E∋y→x} u(y). In Theorem 6.2 part of the conclusion is that this extension is continuous in Ω.

Note also that some condition of the type (a)–(d) is needed: Consider a bounded open set Ω ⊂ R^{n} (unweighted), n ≥ 3, p = 2, and let v be the Green function of Ω with respect to some y ∈ Ω (or v(x) = |x − y|^{2−n}). Then −v is a harmonic function in Ω \ {y} with no quasisuperharmonic extension to Ω.

Let us next observe that unique removability for bounded Q-quasisuperhar-monic functions is the same as for Q-quasisuperharmonic functions bounded from below. Note that in Section 10 we give examples of sets which are nonuniquely removable for p-harmonic functions, but we do not know if there are any nonuniquely removable sets for superharmonic functions (or for Q-quasisuperharmonic functions).

Proposition 6.4. The set E is uniquely removable for Q-quasisuperhar-monic functions bounded from below on Ω\E if and only if it is uniquely removable for bounded Q-quasisuperharmonic functions on Ω \ E.

Proof. The necessity is obvious. As for the sufficiency assume that E is uniquely removable for bounded Q-quasisuperharmonic functions on Ω \ E. Let u be a Q-quasisuperharmonic function bounded from below on Ω\E. Then u_{j} := min{u, j} has a Q-quasisuperharmonic extension U_{j} to Ω. As min{U_{j+1}, U_{j}} = u_{j} in Ω \ E, both U_{j} and min{U_{j+1}, U_{j}} are Q-quasisuperharmonic extensions of u_{j} to Ω. By uniqueness U_{j} = min{U_{j+1}, U_{j}} and thus U_{j+1} ≥ U_{j} in Ω. Let U = lim_{j→∞} U_{j}. Then U = u in Ω and hence U is not identically ∞ in any component of Ω. By Proposition 3.7, U is a Q-quasisuperharmonic extension of u to Ω.

As for the uniqueness, let U be a Q-quasisuperharmonic extension of u to Ω. Then min{U, k} is the unique Q-quasisuperharmonic extension of min{u, k} to Ω. It follows that U is unique.

Proof of Theorem 6.3. The uniqueness follows from the observation above, we let

U(x) = ess lim inf_{Ω\E∋y→x} u(y), x ∈ Ω.

Assume first that (c) holds. It follows that u ∈ N_{loc}^{1,p}(Ω \ E), moreover g_{u} is a p-weak upper gradient of U in Ω \ E. Let Γ_{E} be the set of curves passing through E. By Lemma 3.6 in Shanmugalingam [30], Mod_{p}(Γ_{E}) = 0, and hence g_{u} is a p-weak upper gradient of U in Ω. Let B ⋐ Ω. Since μ(E) = 0, we see that ∥U∥_{N^{1,p}(B)} = ∥u∥_{N^{1,p}(B\E)} < ∞, and thus U ∈ N_{loc}^{1,p}(Ω). Moreover, if u ∈ N^{1,p}(Ω \ E), then U ∈ N^{1,p}(Ω).

We shall now show that U is a Q-quasisuperminimizer in Ω. Let φ ∈ N_{0}^{1,p}(Ω) be nonnegative, and let φ′ = φχ_{Ω\E}. Then φ′ ∈ N_{0}^{1,p}(Ω \ E). Since U is a Q-quasisuperminimizer in Ω \ E, we see that (using characterization (c) in Proposition 3.2),

∫_{φ≠0} g_{U}^{p} dμ = ∫_{φ′≠0} g_{U}^{p} dμ ≤ Q ∫_{φ′≠0} g_{U+φ′}^{p} dμ = Q ∫_{φ≠0} g_{U+φ}^{p} dμ.

Thus U is a Q-quasisuperminimizer in Ω, and since U is lower semicontinuously regularized, U is Q-quasisuperharmonic in Ω.

That (a) ⇒ (c) and (b) ⇒ (c) hold are clear.

Let us consider the following condition:

(e) u is bounded. We next want to show that (e) ⇒ (c). Assume therefore that (e) holds. Let

2B ⋐ Ω and η ∈ Lip_{c}(2B) be such that 0 ≤ η ≤ 1 in Ω and η = 1 in B. Since E ∩ \overline{2B} is compact and has zero capacity, there exists a sequence ψ_{j} ∈ Lip_{c}(X) such that ∥ψ_{j}∥_{N^{1,p}(X)} → 0, 0 ≤ ψ_{j} ≤ 1 in X and ψ_{j} = 1 on E ∩ \overline{2B}, see Theorem 1.1 in Kallunki–Shanmugalingam [17] and Proposition 4.4 in Heinonen– Koskela [13]. We may also assume that ψ_{j} → 0 μ-a.e. Let η_{j} = η(1−2ψ_{j})_{+} . Then η_{j} ∈ Lip_{c}(Ω \ E), 0 ≤ η_{j} ≤ 1 and η_{j} → η in N^{1,p}(X). Let v = −u. Without loss of generality \frac{1}{2} ≤ v ≤ 1. Using the Caccioppoli type estimate (Theorem 4.1) we see that

∫_{B\E} g_{u}^{p}η_{j}^{p} dμ ≤ ∫_{Ω\E} g_{v}^{p}η_{j}^{p} dμ ≤ c ∫_{Ω\E} v^{p}g_{η_{j}}^{p} dμ ≤ c ∫_{Ω\E} g_{η_{j}}^{p} dμ.

By Fatou’s lemma

∫_{B\E} g_{u}^{p} dμ ≤ lim inf_{j→∞} ∫_{B\E} g_{u}^{p}η_{j}^{p} dμ ≤ c lim inf_{j→∞} ∫_{Ω\E} g_{η_{j}}^{p} dμ = c ∫_{Ω\E} g_{η}^{p} dμ < ∞.

Since u is bounded we conclude that u ∈ N^{1,p}(B \ E), and thus (c) holds.

Assume now that (d) holds. As we have shown that E is uniquely removable for bounded Q-quasisuperharmonic functions on Ω \ E, it follows from Proposition 6.4 that u has a Q-quasisuperharmonic extension to Ω, which must equal U.

In order to obtain Theorem 6.2 from Theorem 6.3 we formulate the following result. In view of Problem 9.2 we make it more general than what is actually needed in this section.

Proposition 6.5. Assume that μ(E) = 0 and that u is a Q-quasiharmonic function in Ω \ E, which has a Q-quasisuperharmonic extension U and a Qquasisubharmonic extension V to Ω. Then U = V is both unique and Qquasiharmonic in Ω. If u is bounded, then U is also bounded, and if u ∈ N^{1,p}(Ω \ E), then U ∈ N^{1,p}(Ω).

Here V is Q-quasisubharmonic if −V is Q-quasisuperharmonic.

Proof. Since U is lower semicontinuously regularized and μ(E) = 0, we have

U(x) = ess lim inf_{Ω∋y→x} U(y) = ess lim inf_{Ω\E∋y→x} u(y), x ∈ Ω.

Thus U is unique. Moreover if u is bounded, then U is bounded.

Let B ⋐ Ω. As V is upper semicontinuous in Ω it is bounded from above in B. It follows that u is bounded from above in B \ E, and hence also U is bounded from above in B. By Theorem 7.3 in Kinnunen–Martio [23], U is a Q-quasisuperminimizer in Ω and U ∈ N_{loc}^{1,p}(Ω). Similarly V ∈ N_{loc}^{1,p}(Ω) is a Q-quasisubminimizer in Ω. Since U = V μ-a.e. in Ω, Corollary 3.3 in Shanmugalingam [30] shows that U = V q.e. in Ω. Thus U is also a Q-quasisubminimizer in Ω. Hence U is a Q-quasiminimizer in Ω, and there exists a Q-quasiharmonic function W such that W = U = u μ-a.e. in Ω. Since both W and u are continuous in Ω \ E, they coincide in Ω \ E. By uniqueness, W = U = V in Ω.

Since Ω\E is open, the minimal p-weak upper gradient g_{U} of U in Ω is also minimal as a p-weak upper gradient in Ω\E. Hence ∥U∥_{N^{1,p}(Ω)} = ∥u∥_{N^{1,p}(Ω\E)} .

Proof of Theorem 6.2. This now follows directly by combining Theorem 6.3 and Proposition 6.5.

We end this section by observing that removability for quasi(super)harmonic functions is the same as removability for quasi(super)minimizers. By saying that removability is unique for Q-quasi(super)minimizers we mean that the extensions are unique up to capacity zero.

Proposition 6.6. The set E is (uniquely) removable for bounded Q-quasi- (super)minimizers in Ω \ E if and only if E is (uniquely) removable for bounded Q-quasi(super)harmonic functions in Ω \ E.

This result can be combined with Proposition 6.4. There is also a similar result showing that removability for Q-quasi(super)minimizers in N^{1,p} is the same as removability for Q-quasi(super)harmonic functions in N^{1,p}.

Proof. Assume first that E is removable for bounded Q-quasi(super)mini-mizers in Ω \ E, and that u is a bounded Q-quasi(super)harmonic function in Ω \ E. Then u is a Q-quasi(super)minimizer in Ω \ E, and has a bounded Qquasi(super)minimizer extension U to Ω. Let U^{∗} be the lower semicontinuous regularization of U, a Q-quasi(super)harmonic function in Ω. Since both U^{∗} and u are lower semicontinuously regularized in Ω \ E and U^{∗} = u q.e. in Ω \ E, we see that U^{∗} = u everywhere in Ω \ E.

Conversely, assume that E is removable for bounded Q-quasi(super)harmonic functions in Ω \ E, and that u is a bounded Q-quasi(super)minimizer in Ω \ E. Let u^{∗} be the lower semicontinuous regularization of u in Ω \ E. Then u^{∗} is Q-quasi(super)harmonic in Ω \ E and has a bounded Q-quasi(super)harmonic extension U^{∗} to Ω. Thus

U = \{ \begin{matrix}U^{∗}, & in E, \\ u, & in Ω \ E, \end{matrix}

is a Q-quasi(super)minimizer extending u to Ω.

If the removability is not unique in one case, then the constructions above can be used to obtain two extensions which are different on a set of positive capacity from which the nonuniqueness of the other case follows.

7. Nonremovability

Let from now on K ⊂ Ω be compact, and recall that E ⊊ Ω is assumed to be relatively closed and such that no component of Ω is completely contained in E.

For unbounded Ω the Dirichlet problem needs to be further studied (with or without a boundary condition at ∞). Without such a theory we are able to give one nonremovability result which holds for unbounded domains.

Proposition 7.1. Assume that for some component G ⊂ Ω the set G \ K is disconnected. Then there is a bounded p-harmonic function in Ω \ K with no quasisuperharmonic extension to Ω.

Note that by Example 10.2 there are relatively closed removable sets E ⊂ Ω for which Ω is connected and Ω \ E is disconnected.

Proof. Let u ≡ 1 in one component of G \ K and 0 otherwise. Then u is a bounded p-harmonic function in Ω\K. Assume that it has a quasisuperharmonic extension U to Ω. Since U is lower semicontinuous there is x ∈ G ∩ K such that U(x) = inf_{G∩K} U. Thus

inf_{G} U = min{U(x), inf_{G\K} u} = min{U(x), 0}

is attained at some point in G. By the minimum principle, U is constant in G. But this contradicts the fact that u is nonconstant in G \ K.

In the rest of this section we will consider bounded open sets Ω.

Proposition 7.2. Assume that Ω is bounded, C_{p}(X \ Ω) > 0 and μ(E) = 0 < C_{p}(K∪E). Then there is a bounded p-harmonic function in N^{1,p}(Ω\(K∪E)) with no quasisuperharmonic extension to Ω.

When we just consider quasiharmonic extensions we can be slightly more general.

Proposition 7.3. Assume that Ω is bounded, C_{p}(X \ Ω) > 0, E has empty interior and C_{p}(K ∪ E) > 0. Then there is a bounded p-harmonic function in N^{1,p}(Ω \ (K ∪ E)) with no quasiharmonic extension to Ω.

Note that in neither Proposition 7.2 nor 7.3 the requirement C_{p}(X \ Ω) > 0 can be omitted, not even when E = ∅, see Example 9.3.

In the case when Ω is bounded and C_{p}(X \ Ω) = 0 we have the following result. (Observe that in this case X = \overline{Ω} is bounded and Ω is connected.)

Proposition 7.4. Assume that Ω is bounded, C_{p}(X\Ω) = 0 and C_{p}(E) > 0. Assume that either

(a) the capacity of Ω ∩ ∂E is not concentrated to one point, i.e. that for every x ∈ Ω ∩ ∂E we have C_{p}((Ω ∩ ∂E) \ {x}) > 0, or

(b) Ω \ E is disconnected.

Then there exists a bounded p-harmonic function in N^{1,p}(Ω\E) with no bounded quasisuperharmonic extension to Ω, nor any quasisuperharmonic extension in N^{1,p}(Ω).

When the (positive) capacity of E is concentrated to one point and Ω \ E is connected, it is possible both to have removability and nonremovability for quasiharmonic functions, see Examples 7.7 and 9.3.

When singleton sets have zero capacity we can say more.

Proposition 7.5. Assume that Ω is bounded and C_{p}(E) > 0. If C_{p}({x}) = 0 for all x ∈ Ω ∩ ∂E, then there is a bounded p-harmonic function in N^{1,p}(Ω \ E) that has no quasiharmonic extension to Ω.

To prove these propositions we need the following lemma.

Lemma 7.6 (Björn–Björn–Shanmugalingam [10, Lemma 8.6]). If Ω is a connected set and C_{p}(E) > 0, then C_{p}(Ω ∩ ∂E) > 0.

A standing assumption in [10] was that Ω is a nonempty bounded connected open set with C_{p}(X \ Ω) > 0. However, the proof of this lemma in [10] does not use the boundedness nor the assumption C_{p}(X \ Ω) > 0, and thus the lemma holds as stated here.

Proof of Proposition 7.2. Since C_{p}(K ∪ E) > 0 we can find a compact set K′ such that K ⊂ K′ ⊂ K ∪ E and C_{p}(K′) > 0. Observe that f(x) = min{dist(K′, x)/ dist(K′, X \ Ω), 1} is a Lipschitz function. Let u = H_{Ω\K′}f. Then u is p-harmonic in Ω \ K′, u = 0 on K′ and u = 1 on X \ Ω. Note that 0 ≤ u ≤ 1.

Suppose there is a quasisuperharmonic function U in Ω such that U = u in the open set Ω \ (K ∪ E). The continuity of u and the lower semicontinuous regularity of U together with the condition μ(E) = 0 imply that U = u in Ω \ K ⊃ Ω \ K′.

Since u ≥ 0, it follows that U ≥ 0 in Ω, by the minimum principle. Lemma 7.6 implies that C_{p}(∂K′) > 0 and by the Kellogg property (Theorem 5.3), there exists a regular point x_{0} ∈ ∂K′, i.e.

lim_{Ω\K′∋y→x_{0}} U(y) = lim_{Ω\K′∋y→x_{0}} u(y) = u(x_{0}) = 0.

By the lower semicontinuity of U, U(x_{0}) = 0. The minimum principle implies that U is constant in the component G ⊂ Ω containing x_{0} . Since C_{p}(X \ G) ≥ C_{p}(X \Ω) > 0, Lemma 7.6 and the Kellogg property (Theorem 5.3) however imply that U is nonconstant in G, a contradiction.

Proof of Proposition 7.3. This proof is almost identical to the proof of Proposition 7.2, we only need to modify the second paragraph a little as follows: Suppose there is a quasiharmonic function U in Ω such that U = u in the open set Ω \ (K ∪ E). The continuity of u and U together with the condition that E has no interior imply that U = u in Ω \ K ⊃ Ω \ K′. The rest of the proof is identical to the proof of Proposition 7.2.

Proof of Proposition 7.4. In case (i), it follows that we can find two disjoint compact sets K_{1}, K_{2} ⊂ Ω∩∂E with positive capacity. Let f = 1 on K_{1} , f = 0 on K_{2} and let u = H_{X\(K_{1}∪K_{2})}f. By the Kellogg property there are regular points x_{1} ∈ ∂K_{1} and x_{2} ∈ ∂K_{2} . It follows that u is nonconstant.

In case (ii), we let u ≡ 1 in one component of Ω \ E and u ≡ 0 in all other components of Ω \ E.

Thus in both cases we have a nonconstant p-harmonic function u in Ω\E. Assume that u has a quasisuperharmonic extension U to Ω, which is either bounded or belongs to N^{1,p}(Ω). Then, by Theorem 6.3, there is a quasisuperharmonic function V on X which is an extension of U and hence of u. By Corollary 3.5, V is constant, which contradicts the fact that u is nonconstant.

Proof of Proposition 7.5. There is a component G of Ω such that C_{p}(G∩E) > 0. Since by Lemma 7.6 we have C_{p}(G ∩ ∂E) > 0, there exists τ > 0 so that C_{p}(\overline{G}_{τ} ∩ ∂E) > 0, where G_{τ} := \{x ∈ G : dist(x, X \ G) > τ\} (if X = G we set G_{τ} = G).

By the Kellogg property (Theorem 5.3) and by the fact that finite subsets of G ∩ ∂E have zero capacity, there exists a sequence \{x_{n}\}_{n=1}^{∞} of points in \overline{G}_{τ} ∩ ∂E that are regular for the open set G \ E. Since \overline{G}_{τ} is compact, without loss of generality we may assume that this sequence converges to a point x_{∞} ∈ G ∩ ∂E, has no other limit points, and moreover consists of distinct points. For each x_{n} in this sequence, let B_{n} = B(x_{n}, r_{n}) be a ball so that \overline{B}_{n} ⊂ G. We can also choose the balls B_{n} to be pairwise disjoint.

It follows from Theorem 1.1 in Kallunki–Shanmugalingam [17] and Proposition 4.4 in Heinonen–Koskela [13], that we can find φ_{n} ∈ Lip_{c}(B_{n}) so that φ_{n}(x_{n}) = 1, 0 ≤ φ_{n} ≤ 1 and ∥g_{φ_{n}}∥_{L^{p}} < 2^{−j}. Let

Φ(x) = ∑_{n=1}^{∞} φ_{2n}(x).

It is easy to see that Φ ∈ N^{1,p}(X) is a bounded lower semicontinuous function which is continuous apart from at x_{∞}. Let u = H_{Ω\E}Φ. Clearly u is a bounded p-harmonic function in N^{1,p}(Ω \ E). We will show that u has no quasiharmonic extension to Ω.

Since Φ is continuous at x_{n}, we see by Corollary 7.2 in Björn–Björn–Shanmugalingam [10] that

lim_{Ω\E∋y→x_{n}} u(y) = Φ(x_{n}).

Note that Φ(x_{n}) = 1 if n is even and Φ(x_{n}) = 0 if n is odd. Hence as x_{∞} is the limit point of the sequence {x_{n}}_{n=1}^{∞}, we obtain a sequence {y_{n}}_{n=1}^{∞} in Ω \ E which converges to x_{∞} and is such that u(y_{n}) ≥ \frac{3}{4} if n is even and u(y_{n}) ≤ \frac{1}{4} if n is odd. Thus u|_{Ω\E} has no continuous extension to the point x_{∞} ∈ Ω ∩ ∂E. Since quasiharmonic functions are continuous, u has no quasiharmonic extension to Ω.

Example 7.7. Let X ⊂ R^{2} be the unit circle with the Euclidean distance and the one-dimensional Lebesgue measure. Let K = {(1, 0)}. Observe that C_{p}(K) > 0. By considering (1, 0) as two points we can find a nonconstant pharmonic function u in X \ K such that

lim_{{(x,y)→(1,0) \atop y<0}} u(x, y) = 0 and lim_{{(x,y)→(1,0) \atop y>0}} u(x, y) = 1.

Since quasisuperharmonic functions on X are constant, by Corollary 3.5, u has no quasisuperharmonic extension to X, and K is not removable.

Note that in this case no (nonempty) set is removable.

8. Characterizations of removability

Combining Theorems 6.2, 6.3 and Proposition 7.2, gives the following consequence. (Note that the condition C_{p}(X \ Ω) is essential, see Example 9.3.)

Corollary 8.1. Assume that Ω is bounded and C_{p}(X \ Ω) > 0. Then the following are equivalent:

(a) C_{p}(K) = 0;

(b) K is removable for bounded Q-quasiharmonic functions in Ω \ K;

(c) K is removable for Q-quasiharmonic functions in N^{1,p}(Ω \ K);

(d) K is removable for bounded Q-quasisuperharmonic functions in Ω \ K;

(e) K is removable for Q-quasisuperharmonic functions in N^{1,p}(Ω \ K).

The following characterization is an immediate consequence of Theorem 6.2 and Proposition 7.5.

Corollary 8.2. Assume that Ω is bounded and that C_{p}({x}) = 0 for each x ∈ Ω ∩ ∂E. Then the following are equivalent:

(a) C_{p}(E) = 0;

(b) E is removable for bounded Q-quasiharmonic functions in Ω \ E;

(c) E is removable for Q-quasiharmonic functions in N^{1,p}(Ω \ E).

(Recall that if some component of Ω is contained in E, then E is not considered removable, by Definition 6.1.)

The following is a consequence of Theorems 6.2, 6.3 and Propositions 7.2 and 7.3.

Corollary 8.3. Assume that Ω is bounded, C_{p}(X \Ω) > 0 and E has empty interior. Then the following are equivalent:

(a) C_{p}(K ∪ E) = 0;

(b) K ∪ E is removable for bounded Q-quasiharmonic functions in Ω \ (K ∪ E);

(c) K ∪ E is removable for Q-quasiharmonic functions in N^{1,p}(Ω \ (K ∪ E)).

If moreover μ(E) = 0, then also the following statements are equivalent to those above:

(d) K ∪ E is removable for bounded Q-quasisuperharmonic functions in Ω\ (K ∪ E);

(e) K ∪E is removable for Q-quasisuperharmonic functions in N^{1,p}(Ω\(K ∪E)).

9. Removable sets with positive capacity

In this section we construct some examples of sets with positive capacity which are removable for bounded p-harmonic functions. Example 9.4 is a little simpler than our first example, but we prefer to start with the following example.

Example 9.1. Let X = R (equipped with the Euclidean distance and the one-dimensional Lebesgue measure). In this one-dimensional setting it is easy to see that p-harmonic functions are exactly the linear functions, i.e. of the form x ↦ ax + b, moreover this is true simultaneously for all p. (Observe that on bounded sets all p-harmonic functions are automatically bounded in this case.) Let Ω = (0, 2) and E = [1, 2). Note that C_{p}(E) > 0 and even μ(E) > 0. Thus if f: Ω \ E → R is a p-harmonic function it is linear and thus directly extends to a p-harmonic function in Ω, i.e. E is removable for p-harmonic functions. Note that the extension is unique.

Let us now look at bounded superharmonic functions. In this setting superharmonic functions are exactly the concave functions (again simultaneously for all p). Thus f(x) = \sqrt{1 − x^{2}} is a bounded superharmonic function in Ω \ E, and since lim_{x→1−} f′(x) = −∞ we cannot find a superharmonic (i.e. concave) extension of f to all of Ω. Thus E is not removable for bounded superharmonic functions.

This example shows that the sets removable for bounded p-harmonic functions do not coincide with the sets removable for bounded superharmonic functions. A natural question to ask is the following question.

Problem 9.2. If E is removable for bounded superharmonic functions in Ω \ E, does it then follow that E is removable for bounded p-harmonic functions in Ω \ E?

If μ(E) = 0, then this follows from Proposition 6.5. Otherwise, if f is a bounded p-harmonic function in Ω \ E, then it, of course, has a subharmonic extension and a superharmonic extension to all of Ω, but it is not clear whether these can always be made to coincide.

In the special case when Ω = X is bounded, we can find a compact set with positive capacity which is removable for quasiharmonic functions.

Example 9.3. Let X = Ω = [0, 1] with the Euclidean distance and the one-dimensional Lebesgue measure. Let K = {1}. Observe that C_{p}(K) > 0. Let u be a quasiharmonic function in X \ K. Assume that u(a) = u(b) for some 0 ≤ a < b < 1. Then in order to quasiminimize in (a, b), u must be constant in [a, b]. It follows that u is a continuous monotone function. Without loss of generality, assume that u is nondecreasing. Then u(0) = inf_{X\K} u, and by the minimum principle, u is constant, and thus extendible to all of X as a p-harmonic function. Thus K is removable for Q-quasiharmonic functions in X \ K.

Superharmonic functions in X \ K are exactly concave nonincreasing functions, whereas all quasisuperharmonic functions in X are constant, by Corollary 3.5. Thus K is not removable for bounded Q-quasisuperharmonic functions in X \ K.

In this case the sets E removable for bounded Q-quasiharmonic functions in X \ E are all the sets [0, a] and [a, 1] for 0 ≤ a ≤ 1, except for X itself.

Example 9.4. Let X = [0, ∞) with the Euclidean distance and the onedimensional Lebesgue measure. Let Ω = [0, 1). Then by the arguments in the previous example, a relatively closed set E ⊂ Ω is removable for bounded Qquasiharmonic functions in Ω \ E if and only if E = [a, 1) for some 0 < a < 1.

10. Nonunique removability

Next we will give an example of a set which is removable for p-harmonic functions, but in which the extensions are not unique. We will consider p-harmonic functions on a graph. Such p-harmonic functions were considered by Holopainen– Soardi [15], [16] and Shanmugalingam [32]. For the reader’s convenience we give a brief explanation of how p-harmonic functions are defined on graphs.

Let 𝒢 = (𝒱 , ℰ) be a connected finite or infinite graph, where 𝒱 stands for the set of vertices and ℰ the set of edges. If x and y are endpoints of an edge we say that they are neighbours and write x ∼ y. Consider an edge as a geodesic open ray of length 1 between its endpoints, and let X = 𝒱 ∪⋃_{e∈ℰ} e be the metric graph equipped with the one-dimensional Hausdorff measure μ.

Let Ω ⋐ X be a domain and assume for simplicity that ∂Ω ⊂ 𝒱 . Then u is a p-harmonic function in Ω if and only if it is linear on each edge in Ω and satisfies

(10.1)  ∑_{y∼x} |u(y) − u(x)|^{p−2}(u(y) − u(x)) = 0 for all x ∈ 𝒱 ∩ Ω.

Example 10.1. Consider the graph 𝒢 = ({1, 2, 3, 4}, {(1, 2), (1, 3), (1, 4)}), let X be the corresponding metric graph, Ω = X \ {2, 3, 4} and E = (1, 3) ∪ (1, 4) ∪ {1}, i.e. Ω \ E is just the open edge (1, 2).

A p-harmonic function u in Ω \ E is linear and thus can be described by giving its boundary values u(1) and u(2). Let now

\begin{aligned}U_{1}(1) & = U_{2}(1) = u(1), \\ U_{1}(2) & = U_{2}(2) = u(2), \\ U_{1}(3) & = U_{2}(4) = u(1), \\ U_{1}(4) & = U_{2}(3) = 2u(1) − u(2) \end{aligned}

and let U_{1} and U_{2} be linear on the edges. It is then easy to check that U_{1} and U_{2} are p-harmonic in Ω (simultaneously for all p). Thus E is removable for p-harmonic functions in Ω \ E (which are automatically bounded). However, the extensions are not unique in this case.

Finally note that the same argument as in Example 9.1 shows that E is not removable for bounded superharmonic functions.

In the next example we show that even if E disconnects Ω it is possible to have removability.

Example 10.2. Consider the graph

𝒢 = ({1, …, 6}, {(1, 2), (1, 3), (1, 4), (4, 5), (4, 6)}),

let X be the corresponding metric graph, Ω = X \ {2, 3, 5, 6} and

E = {1, 4} ∪ (1, 3) ∪ (1, 4) ∪ (4, 6).

Then Ω \ E is disconnected and consists of the two edges (1, 2) and (4, 5).

Let u be any p-harmonic function in Ω \ E, which can be described by its boundary values. We want to extend it to a p-harmonic function U in Ω. Again U is described by its boundary values. Apart from being linear at each edge we only need to require U to satisfy (10.1) at the internal edges, and this holds if and only if

\begin{aligned}0 & = φ(U(2) − U(1)) + φ(U(3) − U(1)) + φ(U(4) − U(1)) \\ & = φ(U(1) − U(4)) + φ(U(5) − U(4)) + φ(U(6) − U(4)), \end{aligned}

where φ(t) = |t|^{p−2}t. Since φ: R → R is onto and we can freely choose U(3) and U(6) this can always be achieved.

We can modify this construction to get a set E that disconnects Ω and is nonuniquely removable.

Example 10.3. Consider the graph

𝒢 = ({1, …, 7}, {(1, 2), (1, 3), (1, 4), (4, 5), (4, 6), (4, 7)}),

let X be the corresponding metric graph, Ω = X \ {2, 3, 5, 6, 7} and

E = {1, 4} ∪ (1, 3) ∪ (1, 4) ∪ (4, 6) ∪ (4, 7).

(Thus we have added the extra edge (4, 7) to Example 10.2.)

Arguing as in Example 10.2 we see that every p-harmonic function in Ω \ E is extendible to Ω. Moreover, since we now have the freedom to choose both U(6) and U(7) we can actually prescribe one of them arbitrarily, which shows that the removability is nonunique.

11. Properties of removable singularities

Let E = E_{1} , E_{2}, …, be sets of capacity zero, relatively closed in the nonempty open sets Ω = Ω_{1} , Ω_{2}, …, respectively. Then E_{j} is (uniquely) removable for bounded p-harmonic functions in Ω_{j} \ E_{j}. Moreover, we have the following properties: (Let Ω′ be a nonempty open set.)

(a) (Removability is independent of the surrounding set.) If E is relatively closed in Ω′, then E is uniquely removable for bounded p-harmonic functions in Ω′ \ E.

(b) (A subset of a removable set is removable.) If E′ ⊂ E and E′ ⊂ Ω is relatively closed, then E′ is uniquely removable for bounded p-harmonic functions in Ω \ E′.

(c) (A countable union of removable sets is removable.) If E′ := ⋃_{j=1}^{∞} E_{j} ⊂ Ω is relatively closed, then E′ is uniquely removable for bounded p-harmonic functions in Ω \ E′.

(d) (A generalization of (a)–(c).) If E′ ⊂ ⋃_{j=1}^{∞} E_{j} and E′ ⊂ Ω′ is relatively closed, then E′ is uniquely removable for bounded p-harmonic functions in Ω′ \ E′. This follows directly from the countable subadditivity of the capacity together

with Theorem 6.2. The corresponding statements for Q-quasi(super)harmonic functions that are bounded or in N^{1,p} follow in the same way.

For removable singularities with positive capacity the situation is quite different, even when we restrict our attention to compact removable singularities.

Let X = [0, 1] with the Euclidean distance and the one-dimensional Lebesgue measure, as in Example 9.3. Recall that in this case the sets E removable for bounded Q-quasiharmonic functions in X \ E are all the sets [0, a] and [a, 1] for 0 ≤ a ≤ 1, except for X itself.

We get the following counterexamples corresponding to (a)–(c) above: (a′) Let K = {0} and Ω′ = [0, 1), then K is removable for bounded Q-quasi-

harmonic functions in X \K, but not for bounded Q-quasiharmonic functions in Ω′ \ K.

(b′) Let K = [0, \frac{1}{2}] and K′ = \{\frac{1}{2}\} ⊂ K, then K is removable for bounded Q-quasiharmonic functions in X \ K, but K′ is not removable for bounded Q-quasiharmonic functions in X \ K′.

(c′) Let K_{1} = {0} and K_{2} = {1}, then K_{j} is removable for bounded Q-quasi-harmonic functions in X \ K_{j}, j = 1, 2, but K_{1} ∪ K_{2} is not removable for bounded Q-quasiharmonic functions in X \ (K_{1} ∪ K_{2}). There is however one similar property that holds also for removable singular-

ities of positive capacity.

Proposition 11.1. Let E_{1}, E_{2} ⊂ Ω be relatively closed and such that no component of Ω is contained in E_{1} ∪ E_{2} . Assume that E_{1} and E_{2} are (uniquely) removable for bounded Q-quasi(super)harmonic functions in Ω \ E_{1} and Ω \ (E_{1} ∪ E_{2}), respectively. Then E_{1} ∪ E_{2} is (uniquely) removable for bounded Q-quasi(super)harmonic functions in Ω \ (E_{1} ∪ E_{2}).

It is implicitly assumed that Ω′ := (Ω \ E_{1}) ∪ E_{2} is open.

Proof. Let u be a bounded Q-quasi(super)harmonic function in Ω′ \ E_{2} = Ω \ (E_{1} ∪ E_{2}). By assumption there is a bounded Q-quasi(super)harmonic extension U′ to Ω′. Thus U′′ := U′|_{Ω\E_{1}} is a bounded Q-quasi(super)harmonic extension of u to Ω \ E_{1} . Since E_{1} is removable, there exists a bounded Qquasi(super)harmonic extension U of U′′ to Ω. It is immediate that U is a bounded Q-quasi(super)harmonic extension of u to Ω.

In the case when E_{1} and E_{2} are uniquely removable we see that U|_{Ω′} is a Q-quasisuperharmonic extension of u to Ω′, and hence is unique (and equal to U′). Therefore U|_{Ω\E_{1}} is also unique (and equal to U′′). Using that E_{1} is uniquely removable, the uniqueness of U follows.

It is not known if being p-harmonic is a local property, and in particular if p-harmonic functions form sheaves, i.e. p-harmonicity in both Ω_{1} and Ω_{2} implies p-harmonicity in Ω_{1} ∪ Ω_{2} . When p-(super)harmonic functions are defined using Cheeger’s definition, see, e.g., Björn–Björn–Shanmugalingam [10], then the sheaf property is known to hold (this follows using the p-Laplace equation and a partition of unity argument, see J. Björn [11]). This means that, e.g., on graphs and in Euclidean R^{n} the sheaf property holds.

On the other hand Q-quasiharmonic functions do not form sheaves.

Proposition 11.2. Assume that X is such that p-(super)harmonic functions form a sheaf. Assume further that Ω′ is open, E ⊂ Ω′ ⊂ Ω and that E is (uniquely) removable for bounded p-(super)harmonic functions on Ω′ \ E. Then E is (uniquely) removable for bounded p-(super)harmonic functions on Ω \ E.

Recall that Ω \ E is assumed to be open.

Proof. Let u be a bounded p-harmonic function in Ω\E. Then u is bounded and p-harmonic in Ω′ \ E, and hence has a bounded p-harmonic extension U′ to Ω′. Let now

U = \{ \begin{matrix}u, & in Ω \ E, \\ U′, & in Ω′. \end{matrix}

Then U is p-harmonic in Ω′ as well as in Ω \ E, and hence by the sheaf property U is p-harmonic in Ω.

If U′ is unique, then any bounded p-harmonic extension U of u to Ω, must satisfy U|_{Ω′} = U′, and hence U is also unique.

The corresponding result for superharmonic functions is proved similarly.

94 Anders Björn

Corollary 11.3. Assume that X is such that p-(super)harmonic functions form a sheaf. Let E_{1}, E_{2} ⊂ Ω be relatively closed and such that no component of Ω is contained in E_{1} ∪ E_{2} . Assume further that E_{1} and E_{2} are (uniquely) removable for bounded p-(super)harmonic functions on Ω \ (E_{1} ∪ E_{2}). Then E_{1} ∪ E_{2} is (uniquely) removable for bounded p-(super)harmonic functions on Ω \ (E_{1} ∪ E_{2}).

It is implicitly assumed that (Ω \ E_{1}) ∪ E_{2} and (Ω \ E_{2}) ∪ E_{1} are open.

Proof. By Proposition 11.2, E_{1} is (uniquely) removable for bounded p- (super)harmonic functions on Ω \ E_{1} . The result then follows from Proposition 11.1.

The results corresponding to Propositions 11.1, 11.2 and Corollary 11.3 for functions in N^{1,p} can be proved analogously. Propositions 11.1, 11.2 and Corollary 11.3 can also be combined with Proposition 6.4.

The author has constructed examples of removable singularities for various classes of analytic functions with a similar character to (a′)–(c′), see [1], [2] and [6], see also Hejhal [14, Example 1, p. 19].

References

[1] Björn, A.: Properties of removable singularities for Hardy spaces of analytic functions. - J. London Math. Soc. 66, 2002, 651–670.

[2] Björn, A.: Removable singularities for analytic functions in BMO and locally Lipschitz spaces. - Math. Z. 244, 2003, 805–835.

[3] Björn, A.: p-harmonic measures and the Perron method for p-harmonic functions. - In: Future Trends in Geometric Function Theory RNC Workshop Jyväskylä 2003, Rep. Univ. Jyväskylä Dep. Math. Stat. 92, Univ. Jyväskylä, Jyväskylä, 2003, 23–29.

[4] Björn, A.: A weak Kellogg property for quasiminimizers. - Preprint, Linköping, 2004.

[5] Björn, A.: Characterizations of p-superharmonic functions on metric spaces. - Studia Math. 169, 2005, 45–62.

[6] Björn, A.: Removable singularities in weighted Bergman spaces. - Czechoslovak Math. J. (to appear).

[7] Björn, A.: A regularity classification of boundary points for p-harmonic functions and quasiminimizers. - In preparation.

[8] Björn, A., and J. Björn: Boundary regularity for p-harmonic functions and solutions of the obstacle problem. - Preprint, Linköping, 2004.

[9] Björn, A., J. Björn, and N. Shanmugalingam: The Dirichlet problem for p-harmonic functions on metric spaces. - J. Reine Angew. Math. 556, 2003, 173–203.

[10] Björn, A., J. Björn, and N. Shanmugalingam: The Perron method for p-harmonic functions. - J. Differential Equations 195, 2003, 398–429.

[11] Björn, J.: Approximation by regular sets in metric spaces. - Preprint, Linköping, 2004.

[12] Heinonen, J., T. Kilpeläinen, and O. Martio: Nonlinear Potential Theory of Degenerate Elliptic Equations. - Oxford University Press, Oxford, 1993.

[13] Heinonen, J., and P. Koskela: A note on Lipschitz functions, upper gradients, and the Poincaré inequality. - New Zealand J. Math. 28, 1999, 37–42.

[14] Hejhal, D. A.: Classification theory for Hardy classes of analytic functions. - Ann. Acad. Sci. Fenn. Ser. A I Math. 566, 1973, 1–28.

[15] Holopainen, I., and P. Soardi: p-harmonic functions on graphs and manifolds. - Manuscripta Math. 94, 1997, 95–110.

[16] Holopainen, I., and P. Soardi: A strong Liouville theorem for p-harmonic functions on graphs. - Ann. Acad. Sci. Fenn. Math. 22, 1997, 205–226.

[17] Kallunki, S. [Rogovin, S.], and N. Shanmugalingam: Modulus and continuous capacity. - Ann. Acad. Sci. Fenn. Math. 26, 2001, 455–464.

[18] Keith, S.: Modulus and the Poincaré inequality on metric measure spaces. - Math. Z. 245, 2003, 255–292.

[19] Keith, S., and X. Zhong: The Poincaré inequality is an open ended condition. - Preprint, Jyväskylä, 2003.

[20] Kilpeläinen, T., J. Kinnunen, and O. Martio: Sobolev spaces with zero boundary values on metric spaces. - Potential Anal. 12, 2000, 233–247.

[21] Kinnunen, J., and O. Martio: The Sobolev capacity on metric spaces. - Ann. Acad. Sci. Fenn. Math. 21, 1996, 367–382.

[22] Kinnunen, J., and O. Martio: Choquet property for the Sobolev capacity in metric spaces. - In: Proceedings on Analysis and Geometry (Novosibirsk, Akademgorodok, 1999), Sobolev Institute Press, Novosibirsk, 2000, 285–290.

[23] Kinnunen, J., and O. Martio: Potential theory of quasiminimizers. - Ann. Acad. Sci. Fenn. Math. 28, 2003, 459–490.

[24] Kinnunen, J., and N. Shanmugalingam: Regularity of quasi-minimizers on metric spaces. - Manuscripta Math. 105, 2001, 401–423.

[25] Koskela, P., and P. MacManus: Quasiconformal mappings and Sobolev spaces. - Studia Math. 131, 1998, 1–17.

[26] Martio, O.: Counterexamples for unique continuation. - Manuscripta Math. 60, 1988, 21–47.

[27] Maz’ya, V. G.: Removable singularities of bounded solutions of quasilinear elliptic equations of any order. - Zap. Nauchn. Sem. Leningrad. Otdel. Mat. Inst. Steklov. (LOMI) 26, 1972, 116–130 (Russian). English transl.: J. Soviet Math. 3, 1975, 480–492.

[28] Serrin, J.: Local behavior of solutions of quasi-linear equations. - Acta Math. 111, 1964, 247–302.

[29] Serrin, J.: Isolated singularities of solutions of quasi-linear equations. - Acta Math. 113, 1965, 219–239.

[30] Shanmugalingam, N.: Newtonian spaces: An extension of Sobolev spaces to metric measure spaces. - Rev. Mat. Iberoamericana 16, 2000, 243–279.

[31] Shanmugalingam, N.: Harmonic functions on metric spaces. - Illinois J. Math. 45, 2001, 1021–1050.

[32] Shanmugalingam, N.: Some convergence results for p-harmonic functions on metric measure spaces. - Proc. London Math. Soc. 87, 2003, 226–246.

[33] Tolksdorf, P.: Remarks on quasi(sub)minima. - Nonlinear Anal. 10, 1986, 115–120.

Received 29 October 2004